# Controlled Latent Diffusion Models for $3$D Porous Media Reconstruction

### v2

@@ Abstract
Three-dimensional digital reconstruction of porous media presents a fundamental challenge in geoscience, requiring simultaneous resolution of fine-scale pore structures while capturing representative elementary volumes. We introduce a computational framework that addresses this challenge through latent diffusion models operating within the EDM framework. Our approach reduces dimensionality via a custom Variational Autoencoder trained in binary geological volumes, improving efficiency and also enabling the generation of larger volumes than previously possible with diffusion models. A key innovation is our controlled unconditional sampling methodology, which enhances distribution coverage by first sampling target statistics from their empirical distributions, then generating samples conditioned on these values. Extensive testing on four distinct rock types demonstrates that conditioning on porosity—a readily computable statistic—is sufficient to ensure a consistent representation of multiple complex properties, including permeability, two-point correlation functions, and pore size distributions. The framework achieves better generation quality than pixel-space diffusion while enabling significantly larger volume reconstruction ($256^{3}$ voxels) with substantially reduced computational requirements, establishing a new state-of-the-art for digital rock physics applications.
@@ 1 Introduction
[…]
@@ 3.1 Reconstruction with Unconditional and Controlled Generative Models
[…]
@@ 4.2.1 Autoencoder
The dimension reduction performed by the autoencoder is linear in each dimension, and does not alter the tensorial dimension of the data in order to retain spatial structure in the latent space. That is, for a chosen reduction factor $F$, an input of dimension $[B,C_{in},L,W,H]$ is encoded to an output $[B,C_{l},\frac{L}{F},\frac{W}{F},\frac{H}{F}]$, where $B$ is the batch size, $C_{in}$ the number of input channels, $C_{l}$ the number of latent channels, and $L,W,H$ the spatial dimensions. For a binary image, we have $C_{in}=1$, and following [30] we take $C_{l}=4$. We choose $F=8$ for our reduction factor, as larger $F$ were observed to produce visually perceptible reconstruction errors. The autoencoder architecture is basically the same as in [30], and we also take a very small regularization hyperparameter $\lambda_{\text{KL}}$, setting it to $\lambda_{\text{KL}}=10^{-4}$.
@@ 5.5.2 Controlled unconditional generation
Remarkably, controlling only on porosity is sufficient to essentially eliminate mode collapse in all considered statistics, and we can also recover the correct relation between permeability [^4] See A.5 for details on how we calculate permeability. and porosity [^5] As discussed in A.1, in all samples considered porosity and effective porosity are essentially equal, and thus we use the terms porosity and effective porosity interchangeably., as shown in Figure 18. Controlling also on the TPC, which increases the training time by $2.5$x, improves the generated distribution of some but not all statistics. A concise comparison between all $256^{3}$ models can be seen in Figure 19.
@@ A.2 Two-point correlation function
[…]
@@ A.3 Pore size distribution
    For each pore space $x\in X^{-1}(1)$, we define $B_{0}(x,r_{x})$ to be the ball centered on $x$ with the largest radius such that $B_{0}(x,r_{x})$ only contains pore space. This creates a set =$\mathcal{B}=\{B_{0}(x,r_{x});x\in X^{-1}(1)\}$.
@@
[…]

### v3

@@ Abstract
Note: The final version of this article was published in Computers and Geosciences, Volume 206, January 2026, 106038. DOI: 10.1016/j.cageo.2025.106038. Readers should refer to the published version for the most up-to-date content. Three-dimensional digital reconstruction of porous media presents a fundamental challenge in geoscience, requiring simultaneous resolution of fine-scale pore structures while capturing representative elementary volumes. We introduce a computational framework that addresses this challenge through latent diffusion models operating within the EDM framework. Our approach reduces dimensionality via a custom Variational Autoencoder trained in binary geological volumes, improving efficiency and also enabling the generation of larger volumes than previously possible with diffusion models. A key innovation is our controlled unconditional sampling methodology, which enhances distribution coverage by first sampling target statistics from their empirical distributions, then generating samples conditioned on these values. Extensive testing on four distinct rock types demonstrates that conditioning on porosity—a readily computable statistic—is sufficient to ensure a consistent representation of multiple complex properties, including permeability, two-point correlation functions, and pore size distributions. The framework achieves better generation quality than pixel-space diffusion while enabling significantly larger volume reconstruction ($256^{3}$ voxels) with substantially reduced computational requirements, establishing a new state-of-the-art for digital rock physics applications.
@@ 1 Introduction
We also introduce a novel sampling methodology, termed controlled unconditional sampling, to enhance the coverage of the target distribution. This approach leverages a conditional generative model by first sampling a target statistic $y$ from its empirical distribution $r(y)$ observed in training data, then generating samples conditioned on this value using our trained conditional model $p(x|y)$. The marginalization over $y$, given by $\int p(x|y)r(y)dy$, approximates the desired unconditional distribution $p(x)$, ensuring comprehensive coverage provided that the conditional model accurately captures underlying relationships.
@@ 3.1 Reconstruction with Unconditional and Controlled Generative Models
$p(\mathbf{x})=\int p(\mathbf{x}\mid\mathbf{y})\,r(\mathbf{y})\,d\mathbf{y}\approx q(\mathbf{x}).$ (1)
@@ 4.2.1 Autoencoder
The dimension reduction performed by the autoencoder is linear in each dimension, and does not alter the tensorial dimension of the data in order to retain spatial structure in the latent space. That is, for a chosen reduction factor $F$, an input of dimension $B,C_{in},L,W,H]$ is encoded to an output $B,C_{l},\frac{L}{F},\frac{W}{F},\frac{H}{F}]$, where $B$ is the batch size, $C_{in}$ the number of input channels, $C_{l}$ the number of latent channels, and $L,W,H$ the spatial dimensions. For a binary image, we have $C_{in}=1$, and following [30] we take $C_{l}=4$. We choose $F=8$ for our reduction factor, as larger $F$ were observed to produce visually perceptible reconstruction errors. The autoencoder architecture is basically the same as in [30], and we also take a very small regularization hyperparameter $\lambda_{\text{KL}}$, setting it to $\lambda_{\text{KL}}=10^{-4}$.
@@ 5.5.2 Controlled unconditional generation
Remarkably, controlling only on porosity is sufficient to essentially eliminate mode collapse in all considered statistics, and we can also recover the correct relation between permeability [^5] See A.5 for details on how we calculate permeability. and porosity [^6] As discussed in A.1, in all samples considered porosity and effective porosity are essentially equal, and thus we use the terms porosity and effective porosity interchangeably., as shown in Figure 18. Controlling also on the TPC, which increases the training time by 2.52.5x, improves the generated distribution of some but not all statistics. A concise comparison between all $256^{3}$ models can be seen in Figure 19.
@@ A.2 Two-point correlation function
$\begin{split}&\rho_{f}(y_{1},y_{2};X_{f}):=\int_{[0,L]^{2}}X_{f}(x_{1}+y_{1},x_{2}+y_{2})X_{f}(x_{1},x_{2})I(x_{1}+y_{1},x_{2}+y_{2})dx_{1}dx_{2}\\ &I(z_{1},z_{2}):=\begin{cases}1&(z_{1},z_{2})\in[0,L]^{2}\\ 0&\text{ otherwise}.\end{cases}\end{split}$ (9)
@@ A.3 Pore size distribution
    For each pore space $x\in X^{-1}(1)$, we define $B_{0}(x,r_{x})$ to be the ball centered on $x$ with the largest radius such that $B_{0}(x,r_{x})$ only contains pore space. This creates a set $\mathcal{B}=\{B_{0}(x,r_{x});x\in X^{-1}(1)\}$.
@@
This defines a function $\operatorname{psd}:X^{-1}(1)\to\mathbb{R}^{+}$ that associates each pore space with a pore size, which we call the pore size distribution (PSD). The mean pore size is given by the mean $\frac{1}{\operatorname{vol}(X^{-1}(1))}\int_{X^{-1}(1)}\operatorname{psd}(x)dx$.